OpenStudy (anonymous):

x= y^2+10y-6, please solve for y

OpenStudy (kc_kennylau):

y^2+10y-(6+x)=0 y=\(\dfrac{-10\pm\sqrt{100-24-4x}}{2}\)

OpenStudy (anonymous):

are you sure

OpenStudy (kc_kennylau):

yes

OpenStudy (anonymous):

x=(y+5)^2-31 find y

OpenStudy (kc_kennylau):

x=(y+5)^2-31 x+31=(y+5)^2 sqrt(x+31)=y+5 sqrt(x+31)-5=y P.S. and i AM sure.

OpenStudy (anonymous):

thx

OpenStudy (kc_kennylau):

no problem